A ball is thrown with a speed of 100m/s attains a height of 150m. calculate the time of flight.
1
Expert's answer
2019-03-22T10:32:30-0400

The time of flight is


"t=2\sqrt{\frac{2H}{g}}"

"t=2\sqrt{\frac{2(150)}{9.8}}=11 s"